A manufacturer of consumer electronic equipment makes full use not only of statistical process control, but also of automated testing equipment that efficiently tests all completed products. Data from the testing equation show that finished produces have only 3.0 defects per million opportunities.

Because the manufacturing quality is so high, the process of writing up orders is the major source of quality problems: the defect rate there is 9000 per million opportunities. The manufacturer process about 600 orders per month.

a) What is ?¯ for the order‑writing process? (Enter your answer rounded to three decimal places.)

?¯=

How many defective orders do you expect to see in a month? (Enter your answer rounded to one decimal place.)

Defective Orders=

(b) What is the center line and control limits for ?p chart for plotting monthly proportions of defective orders? Because p^≥0 , if ???<0,  enter 0 . (Enter your answers rounded to four decimal places.)

??=

???=

???=

What is the smallest number of bad orders in a month that will result in a point above the upper control limit? (Enter your answer rounded to the nearest whole number.)

Bad Orders=
Answer #1

a)

phat=9000*10-6 =0.009

defective orders do you expect to see in a month =600*0.009=5.4

b)

CL =0.0090

LCL min(0.009-3*sqrt(0.009*0.991/600),0)=0

UCL =0.0206

Bad Orders= 13